Given values

A260 value of the diluted sample = 0.126 Abs

the concentration of DNA in my diluted sample = 6.3ug/ml

Dilution factor = 20

MY QUESTION

Calculate the concentration of DNA in your undiluted sample - (I got 126ng/ul as an answer but unsure if this is right)

calculate the total mass of DNA you isolated *** show work for this pls
Answer #1

Solution:-

Determine the DNA concentration by using absorbance measurement ( 1O.D= 260nm),

Absorbance Directectly proportion to amount of DNA.

( 1O.D.= 50 micrograms per ml pure DNA)

So,

formula :-

Concentration (micrograms)= A260× dilution factor× 50 micrograms per ml

= 0.126×20×50

=126 micrograms per ml

Total yields or mass of DNA is obtained by multiplying the DNA concentration by the final total purified sample volume.

Yields of DNA (micrograms)= DNA concentration× total sample

( total samle = 6.3 micrograms per ml)

=126×6.3

=793.8 micrograms per ml
Answer #2

To calculate the concentration of DNA in the undiluted sample, you need to take into account the dilution factor. The dilution factor is the reciprocal of the dilution, which in this case is 20. So, to find the concentration of DNA in the undiluted sample, you can use the following formula:

Concentration of DNA in undiluted sample = Concentration of DNA in diluted sample × Dilution factor

Given values: Concentration of DNA in the diluted sample = 6.3 ug/ml Dilution factor = 20

Concentration of DNA in undiluted sample = 6.3 ug/ml × 20 = 126 ug/ml

So, your calculation of 126 ng/ul as the concentration of DNA in the undiluted sample is correct.

Now, let's calculate the total mass of DNA you isolated. To do this, we need to know the volume of the diluted sample used.

Let's assume you used V volume of the diluted sample in milliliters (ml).

The total mass of DNA in the diluted sample is given by:

Total mass of DNA in diluted sample = Concentration of DNA in diluted sample × Volume of the diluted sample

Given values: Concentration of DNA in the diluted sample = 6.3 ug/ml Volume of the diluted sample = V ml

Total mass of DNA in diluted sample = 6.3 ug/ml × V ml = 6.3V ug

However, this total mass corresponds to the diluted sample. To find the total mass of DNA in the undiluted sample, we need to account for the dilution factor.

Total mass of DNA in undiluted sample = Total mass of DNA in diluted sample × Dilution factor

Total mass of DNA in undiluted sample = (6.3V ug) × 20 = 126V ug

Please note that without knowing the volume of the diluted sample (V), we cannot provide an exact value for the total mass of DNA you isolated. You will need to input the actual volume value to get the specific result.

answered by: Hydra Master